# Time-Scale Coupling Between States and Parameters in Recurrent Neural Networks

### v2

@@ Results.
Interpretation.
The slope $s(\ell)$ provides a quantitative diagnostic of how gates renormalize effective learning rates:
  • $s=0$: no lag dependence, i.e. the effective learning rate remains constant and equal to the nominal $\mu$.
  • $s=1$: exact agreement with the zeroth-order gate-product predictor, i.e. $\tilde{\mu}_{\mathrm{eff}}(h)\propto P_{t,k}$.
  • $s>1$: effective sensitivity decays more steeply with lag than the predictor, shortening the memory horizon and yielding more conservative updates.
  • $0<s<1$: effective sensitivity decays more slowly with lag than the predictor, extending the memory horizon and yielding more permissive updates.
The fact that $s(\ell)$ is neither 0 nor $1$ in practice confirms the presence of first-order correction terms beyond the pure gate product. Importantly, $s(\ell)$ remains relatively stable across training, suggesting that this modulation is a structural property of the dynamics rather than a transient effect of optimization.
Task dependence of the slope.
The fitted log–log slope $s(\ell)$ is not an architecture-only quantity: it depends on how the data and objective shape the Jacobians $J_{j}(\theta_{\ell})$ through the activation regime and the statistics of the gates. Tasks that keep units in a near-linear regime and encourage persistently open gates tend to yield $0<s<1$ (slower-than-predictor decay of the empirical effective learning rate with lag), whereas tasks that drive saturation or stronger filtering typically yield $s>1$ (faster-than-predictor decay). Practically, this means the naive gate-product predictor is renormalized by first-order effects in a task-specific manner. Conditioning the fit on gate-product or activation-slope quantiles leads to consistent shifts in $s$, and longer-horizon tasks generally reduce $s$ by encouraging memory-preserving dynamics. Thus, observed deviations of $s$ from $1$ quantify meaningful, task-induced corrections beyond the zeroth-order gate product and should be interpreted as part of the learned, problem-dependent time-scale interaction.
Figures.
We include, for each model, (left) a representative profile at the final checkpoint with both theory overlays, (middle) the evolution of the fitted slope $s(\ell)$ across training, and (right) the sensitivity heatmap $S_{t,k}$ showing the full temporal structure.
7.2 Anisotropy of gradient propagation: gates vs. optimizer
In recurrent networks, the backpropagated error signal at time $k$ is obtained by transporting the loss gradient at time $t$ through the Jacobian products
$M_{t,k}\;=\;\prod_{j=k+1}^{t}J_{j}\,.$
These operators control not only the magnitude of gradients (how much signal survives with lag) but also their directional structure. Any gradient vector reaching earlier timesteps is linearly transformed by $M_{t,k}$; if the spectrum of $M_{t,k}$ is concentrated in a few large singular values, then error signals from many different losses will be funneled into the same dominant subspace. In this case, gradient flow becomes effectively low-dimensional, biasing parameter updates toward specific search directions. Conversely, if singular values remain comparable, gradients are spread across many directions, leading to more isotropic updates.
We therefore define anisotropy of gradient propagation as the degree of directional concentration induced by the singular spectrum
$\sigma_{1}(M_{t,k})\;\geq\;\cdots\;\geq\;\sigma_{N_{r}}(M_{t,k})\,.$
A rapid decay in $\sigma_{i}$ with $i$ indicates that long-range sensitivities align with a restricted set of directions, whereas a flatter spectrum reflects a more isotropic distribution of gradient energy across state space.
This section quantifies such anisotropy and compares three models: (i) a plain RNN trained with Adam (no gates, adaptive optimizer), (ii) a scalar-gated RNN trained with SGD (15), and (iii) a multi-gated RNN (16) trained with SGD. The central question is whether neuron-wise gates with a fixed global step size induce directional concentration of temporal sensitivities that is comparable to, or stronger than, optimizer-induced preconditioning.
Quantities.
For each probe sequence and checkpoint $\ell$, we form Jacobian products $M_{t,k}(\theta_{\ell})=\prod_{j=k+1}^{t}J_{j}(\theta_{\ell})$ and compute (via svd) its singular values $\sigma_{1}\!\geq\!\cdots\!\geq\!\sigma_{N_{r}}$. We aggregate over all pairs with lag h=(tk)h{=}(t{-}k) to obtain two lag-indexed metrics:
  • Anisotropy Index (AI): $\mathrm{AI}_{r}(h;\ell)=\mathrm{median}_{h}\big{(}\sigma_{1}/\sigma_{r}\big{)}$, which is large when sensitivity mass concentrates along a leading direction. We use r=8r{=}8.
  • Energy concentration (CE): $\mathrm{CE}_{r}(h;\ell)=\mathrm{median}_{h}\big{(}\sum_{i=1}^{r}\sigma_{i}^{2}/\sum_{i}\sigma_{i}^{2}\big{)}$, which lies in $[0,1]$ and is high if the top $r$ modes dominate. We use r=8r{=}8.
Both metrics summarize the directional geometry of temporal credit assignment: AI emphasizes the spread between the first and $r$-th mode; CE measures how much of the total sensitivity energy sits in the top subspace.
Setup.
We train the three models above on two standard sequence problems: (i) the adding task (real-valued, two marked positions), and (ii) a compact copy task (one-hot tokens followed by a delimiter, with a regression proxy target). All models have the same state dimension and are trained to comparable loss (plain RNN uses Adam, gated models use plain SGD; no momentum). At selected iterations $\ell\in\{0,400,800\}$ we run a fixed probe batch, build $J_{j}(\theta_{\ell})$, form products $M_{t,k}$ for lags $h\in\{1,2,4,8,12,16,24,32\}$, and record AI and CE per lag (median and interquartile range).
Results (adding task).
At initialization, the multi-gated model trained with SGD already exhibits pronounced directional concentration for medium and long lags, with a larger anisotropy index (AI $\sigma_{1}/\sigma_{r}$) than the plain RNN with Adam at several lags (Figure 4). As training proceeds, the plain RNN with Adam rapidly increases its anisotropy: by the final checkpoint (iter 800) it attains the highest AI across nearly all lags, whereas the multi-gated model remains consistently anisotropic but below Adam, and the scalar-gated model stays lowest among the three. The energy-concentration metric (CEr) mirrors these trends: CE rises with lag for all models, with Adam ultimately achieving the strongest concentration (largest CE) at the last checkpoint, multi-gate+SGD intermediate, and scalar gate weakest. Taken together, these curves indicate that optimizer-induced preconditioning (Adam) can produce stronger low-dimensional amplification of temporal sensitivities than gates alone on this task; nevertheless, neuron-wise gates with a fixed step size still induce substantial, structured anisotropy without any adaptive optimizer.
Results (copy task).
On the copy task, all three models display growing anisotropy with lag and training. The multi-gated RNN with SGD consistently achieves the highest anisotropy index across checkpoints and lags, indicating that per-neuron gating is particularly effective at biasing temporal dynamics toward dominant low-dimensional directions. The plain RNN trained with Adam does increase its anisotropy compared to initialization, but remains below the multi-gate baseline throughout training. Moreover, in the CEr plots Adam exhibits a distinct drop at larger lags (24 and 32), suggesting that the optimizer-induced concentration weakens when gradient propagation is stretched further in time. The scalar-gated model again provides only modest anisotropy, positioned between plain SGD and multi-gate but substantially below both at long horizons.
Takeaways.
(i) All models exhibit increasing anisotropy with lag, confirming that long-range gradient flow becomes effectively low-dimensional. (ii) Per-neuron gates with a fixed global step size yield the most robust and consistent anisotropy, outperforming Adam across training on the copy task. (iii) Adam can boost anisotropy relative to a plain RNN, but its effect is task-dependent: on the adding task it eventually surpasses the multi-gate baseline, whereas on the copy task it remains weaker and loses concentration at longer lags. (iv) The scalar gate provides some concentration but markedly less than multi-gate. These findings complement the results in Section 7.1: beyond scaling the magnitude of effective learning rates with lag, gates and optimizers also sculpt which directions in state space carry long-range gradient information, with multi-gate architectures proving especially effective for tasks requiring persistent memory.
Figure 4: Early training. Anisotropy is already detectable for multi-gate and grows with lag; the relative ordering across models is preserved throughout training.
Figure 5: Adding task. Multi-gate (SGD) exhibits the strongest directionality (higher AI, higher CE) across lags; scalar gate (SGD) is intermediate; plain RNN (Adam) is lowest. Error bands show interquartile ranges across $(t,k)$ pairs at each lag.
Figure 6: Copy task. Directional focusing with multi-gate is stronger and emerges at smaller lags than on the adding task; scalar gate remains intermediate; plain+Adam concentrates more gradually.
7.3 Broader implications
@@
The simulations of Sections 7.1 and 7.2 reveal two complementary roles of gating and optimizer dynamics. First, gates modulate the magnitude of effective learning rates across lags by controlling how much past information is preserved in the state trajectory. Second, gates also shape the directional structure of gradient propagation, determining which subspaces of the state space remain active over long horizons.
These findings highlight that gating mechanisms do more than just stabilize training: they act as structural priors that bias the geometry of gradient flow. Optimizer adaptivity (e.g. Adam) can induce similar concentration effects, but the results show that this behavior is task-dependent: on the adding problem, Adam eventually matches or exceeds the anisotropy induced by multi-gating, while on the copy task its concentration weakens at longer lags. In contrast, per-neuron gating produces robust anisotropy across tasks. The scalar gate baseline sits between these extremes, offering partial control over anisotropy but without the full flexibility of neuron-wise gating.
@@
Taken together, these results suggest a general design principle: architectures and optimizers jointly determine not only how fast parameters move, but also which temporal and spatial directions dominate learning. This dual role has direct relevance for modern sequence models. In gated RNNs, the anisotropy induced by gates may interact with optimizer preconditioning in subtle ways; in attention-based models, analogous phenomena may arise through value/attention weighting. Our framework provides a lens to study such interactions systematically, pointing toward hybrid strategies where structural gating and optimizer dynamics are co-designed to control both effective learning rates and anisotropy of temporal credit assignment.
8 Conclusions and future directions
We developed a theory of time-scale interaction between state and parameter dynamics in RNNs, showing how gating mechanisms not only control information flow in state-space, but also simultaneously regulate effective learning rates and the anisotropy of gradient propagation. Our simulations confirmed that: (i) per-neuron gates reliably extend the memory horizon and induce strong anisotropy, (ii) scalar gates offer intermediate control, (iii) adaptive optimizers such as Adam can mimic or even surpass gate-induced anisotropy in some tasks, but their behavior remains task-dependent, and (iv) the interplay between gates and optimizers determines both how much and which directions of past information shape parameter updates.
These findings suggest that gates should be understood also as structural priors for temporal geometry, while optimizers act as dynamic reweighting mechanisms. This perspective opens new avenues for co-designing architectures and optimizers to achieve desired trade-offs in memory, stability, and efficiency of sequence learning. Extensions to LSTMs [13] and GRUs [6] and transformers [34] constitute promising directions for future work.
Appendix A Matrix product expansion via the Fréchet derivative formulation
@@
A.1 Fréchet differentiability and the first-order expansion
@@
Definition A.1 (Fréchet differentiability [19, 12]).
@@
where $L_{g}(A,E)$ denotes the Fréchet derivative of $g$ at $A$ in the direction $E$; see [12, Sec. 3.2, Thm. 3.3] for a proof.
A.2 Matrix products with structured perturbations
@@
A.3 Recursive application of the product rule
@@
A.4 First-order expansion
@@
$F(\varepsilon)\;=\;\left(\prod_{j=1}^{n}A_{j}\right)\;+\;\varepsilon\sum_{m=1}^{n}\left(\prod_{j=m+1}^{n}A_{j}\right)\,B_{m}\,\left(\prod_{j=1}^{m-1}A_{j}\right)\;+\;O(\varepsilon^{2}).$ (69)
A.5 Simulations supporting the validity of the first-order approximation

### v3

@@ Results
Takeaways
The fitted log–log slope $s(\ell)$ is never exactly 0 or $1$, confirming two key aspects of our theory. First, since $s(\ell)\neq 0$, the empirical effective learning rate always diverges from the nominal learning rate, validating that gates structurally modulate optimization. Second, deviations from $s(\ell)=1$ indicate the presence of first-order correction terms beyond the pure gate product. These corrections are not mere transient effects of optimization: $s(\ell)$ remains relatively stable across training, suggesting that the modulation is a structural property of the dynamics. At the same time, $s(\ell)$ is shaped by the task and data: problems that keep units in a near-linear regime with open gates yield $0<s(\ell)<1$ (slower-than-predictor decay), whereas tasks driving saturation or stronger filtering tend to produce $s(\ell)>1$ (faster-than-predictor decay). In practice, the naive gate-product predictor is thus renormalized by first-order, task-dependent effects, so that observed deviations of $s(\ell)$ from unity quantify meaningful, problem-induced corrections to the effective learning rate.
VII-B Directional anisotropy: propagation vs. updates
Backpropagated signals in recurrent models are transported by Jacobian products
$M_{t,k}=\prod_{j=k+1}^{t}J_{j},$ (52)
which determine both how much gradient survives with lag (magnitude) and how it is distributed across directions (geometry). A concentrated singular spectrum of $M_{t,k}$ indicates that sensitivities align with a low-dimensional subspace in state space. However, optimization updates follow the realized gradients, which depend on the data and on how the loss projects onto those directions. Therefore we report two complementary measurements:
(A) Propagation anisotropy (from Jacobians). For many $(t,k)$ pairs with lag h=tkh=t-k and a fixed probe batch, we compute the singular values $\sigma_{1}\geq\sigma_{2}\dots$ and summarize them by:
$\mathrm{AI}_{r}(M)=\operatorname{median}_{h}\!\Big{(}\tfrac{\sigma_{1}}{\sigma_{r}}\Big{)},$ (53)
$\mathrm{CE}_{r}(M)=\operatorname{median}_{h}\!\Big{(}\tfrac{\sum_{i=1}^{r}\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}\Big{)}.$ (54)
AI emphasizes spectral spread; CE captures how much energy sits in the top-$r$ subspace. Medians are taken across $(t,k)$ pairs with the same lag $h$.
(B) Parameter update anisotropy (from gradient covariance). We collect per-sample gradients into a matrix $G\in\mathbb{R}^{m\times p}$, where each of the $m$ rows contains the flattened gradient of all $p$ trainable parameters for one probe sequence. To isolate directional geometry, each row is $\ell_{2}$-normalized, columns are mean-centered, near-constant columns are dropped, and a small jitter is added for numerical stability. The singular values of $G$ are proportional to the square roots of the eigenvalues of the gradient covariance matrix $C=G^{\top}G$, so spectral analysis of $G$ directly reveals the anisotropy of parameter update directions. If $\sigma_{1}\geq\sigma_{2}\geq\dots$ are the singular values of $G$, we report
$\mathrm{AI}_{r}(G)=\tfrac{\sigma_{1}}{\sigma_{r}},\qquad\mathrm{CE}_{r}(G)=\tfrac{\sum_{i=1}^{r}\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}.$ (55)
These metrics quantify how strongly parameter updates concentrate into a low-dimensional subspace under the training distribution.
Setup
We compare three models with the same state size: (i) plain RNN trained with Adam (no gates), (ii) scalar-gated RNN (11) trained with SGD, and (iii) multi-gated RNN (12) trained with SGD (per-neuron gates). We evaluate on five canonical sequence tasks: adding, AR(2), delay-sum, moving-average, and NARMA10. For each task and model, we report AI/CE of Jacobian products vs. lag, and AI/CE from the gradient covariance. In all tasks, we use $r=10,m=256$ and train the models for 1200 iterations with no early stopping criterion.
Results
A consistent picture emerges when comparing propagation anisotropy (from Jacobians) with parameter update anisotropy (from gradient covariances):
  • Jacobians. Across all tasks, anisotropy indices (AI) rise steeply with lag and CE10 approaches 1, confirming that long-range transport of error signals is effectively low-dimensional. Strikingly, the plain RNN with Adam often exhibits the largest Jacobian AI at long lags (adding, AR(2), delay-sum, moving-average; cf. Figs. 1215), largely because small singular values collapse at large $h$. For NARMA10, the same tendency holds, though by the final checkpoint the difference between Adam and gated models narrows.
[…]
  • Perturbative contributions. The relative advantage of scalar vs. multi-gate depends on task structure. Multi-gate excels on nonlinear or strongly interacting dynamics (moving-average, NARMA10), whereas the scalar gate is competitive—and in some linear aggregation tasks (AR(2), delay-sum) occasionally stronger. This can be understood in light of the perturbative contributions derived in Equation (35) and (43): even though the leading-order effect of a scalar gate is rank-1, the correction terms vary with lag and can redirect sensitivity into different subspaces. As a result, scalar gating may still induce nontrivial anisotropy, albeit in a more fragile and task-dependent manner than neuron-wise gating.
Takeaways
All models develop low-dimensional propagation at long lags, but this can overstate the directional structure of updates. Jacobian spectra capture what the dynamics could amplify, whereas the gradient covariance reveals what optimization actually excites. Adam’s preconditioning often inflates Jacobian anisotropy by shrinking trailing singular values, yet this does not guarantee alignment between loss gradients and dominant transport directions. In contrast, gating reshapes the state dynamics so that realized gradients concentrate into coherent, low-dimensional subspaces of parameter space. Across all tasks, gated models induce markedly stronger update anisotropy than Adam, with near-total energy in the top-$r$ directions. The relative advantage of per-neuron vs. scalar gating is task-dependent: multi-gate dominates on nonlinear or strongly interacting dynamics, while scalar gating can match or exceed it on linear aggregation tasks. Importantly, even scalar gating introduces perturbative contributions (cf. Eqs. (35), (43)) so that, although $G$ is rank-1 at leading order, distinct lags can excite different directions and produce nontrivial anisotropy. For multi-gate models these perturbative contributions are full-rank, further enriching the spectrum of possible update directions.
VII-C Broader implications
@@
The simulations of Sections VII-A and VII-B reveal two complementary roles of gating and optimizer dynamics. Gates modulate the magnitude of effective learning rates across lags by controlling how much past information is preserved in the state trajectory, and they also shape the directional structure of gradient flow—both through Jacobian products and through the geometry of gradient covariance.
A central insight from Section VII-B is that propagation anisotropy and update anisotropy need not coincide. Jacobian spectra show that all models, including plain RNNs with Adam, become increasingly anisotropic as lag grows, with Adam often the most extreme at long horizons. Yet the gradient covariance analysis tells a different story: across benchmark tasks (adding, AR(2), delay-sum, moving-average, NARMA10), scalar and multi-gated models exhibit much stronger update anisotropy, concentrating gradient energy in a low-dimensional subspace. Adam rarely achieves comparable concentration, even when its Jacobian metrics appear favorable. This divergence underscores that while optimizers can influence what the system could amplify, gates determine what the loss actually excites during training.
@@
Overall, gating mechanisms act as structural priors that bias both the magnitude and geometry of credit assignment. Optimizer adaptivity (e.g. Adam) can increase propagation anisotropy in a task-dependent way, but per-neuron gating consistently produces robust concentration of updates, while scalar gating suffices for simpler linear dynamics. Taken together, these results suggest a design principle: architectures and optimizers jointly determine not only how fast parameters move, but also which temporal and spatial directions dominate learning. This dual role is directly relevant for modern sequence models, where gate-induced alignment between state dynamics and gradient directions may complement or compete with optimizer preconditioning, and where analogous effects in attention-based models arise through selective weighting of past activations. Our framework offers a systematic lens to disentangle these contributions and points toward hybrid strategies in which gating and optimizer dynamics are co-designed to control both effective learning rates and the anisotropy of temporal credit assignment.
VIII Conclusions and future directions
This work developed a unified dynamical systems theory of how gates in recurrent neural networks couple state evolution with parameter updates. We showed that gating mechanisms are not merely devices for controlling information flow in state space: they also act as implicit, data-driven preconditioners that modulate effective learning rates across lags and shape the directional geometry of gradient propagation. Through exact Jacobian derivations and first-order expansions, we made explicit how constant, scalar, and multi-dimensional gates induce adaptive step-size behavior, introduce anisotropy in parameter updates, and connect formally to classical optimization strategies such as learning-rate schedules, momentum, and Adam. These results position gating as a structural mechanism that embeds temporal geometry directly into the dynamics of learning, independent of any external optimizer.
The empirical simulations corroborate these theoretical insights across diverse synthetic sequence tasks. They confirm that gates systematically reshape both the magnitude and geometry of temporal credit assignment, producing low-dimensional update structure akin to—and in some cases stronger than—adaptive optimization methods.
Future directions include extending this framework to LSTMs [33], GRUs [34], and transformers [35], as well as exploring hybrid strategies where gating and optimizer dynamics are deliberately co-tuned to balance stability, efficiency, and representational power.
Supplementary material
VIII-A Matrix product expansion via the Fréchet derivative formulation
@@
VIII-A1 Fréchet differentiability and the first-order expansion
@@
Definition VIII.1 (Fréchet differentiability [37, 36]).
@@
where $L_{g}(A,E)$ denotes the Fréchet derivative of $g$ at $A$ in the direction $E$; see [36, Sec. 3.2, Thm. 3.3] for a proof.
VIII-A2 Matrix products with structured perturbations
@@
VIII-A3 Recursive application of the product rule
@@
VIII-A4 First-order expansion
@@
$F(\varepsilon)\;=\;\underbrace{\left(\prod_{j=1}^{n}A_{j}\right)}_{\text{dominant dynamics}}\;+\;\underbrace{\varepsilon\sum_{m=1}^{n}\left(\prod_{j=m+1}^{n}A_{j}\right)\,B_{m}\,\left(\prod_{j=1}^{m-1}A_{j}\right)}_{\text{perturbative correction terms}}\;+\;O(\varepsilon^{2}).$ (66)
VIII-A5 Simulations supporting the validity of the first-order approximation
@@ VIII-B Adam
VIII-C Figures of the simulations
Figure 12: Adding task. Left/middle: propagation anisotropy (AI, CE) vs. lag. Bottom: update anisotropy from gradient covariance (higher is more concentrated).
Figure 13: AR(2). Propagation is highly anisotropic for all models; updates concentrate much more with gates.
Figure 14: Delay-sum. Update anisotropy is extreme for scalar/multi gates; Adam remains much flatter.
Figure 15: Moving-average. Multi-gate shows the strongest update concentration; scalar is a close second.
Figure 16: NARMA10. The gap between gated and Adam models is largest in update anisotropy.